### v2

# Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning

Abstract
Prompt tuning methods have shown comparable performance to general training methods as parameter-efficient fine-tuning (PEFT) methods in various natural language understanding tasks. However, existing prompt tuning methods still utilize the entire model architecture even when solving a specific task, which prevents them from accelerating inference speed during the application procedure. In this paper, we propose a novel prompt tuning framework called Skeleton to efficiently utilize a language model in terms of memory and time complexity for solving various tasks, retaining only task-relevant neurons by using an explainability method. From our framework, we can efficiently solve various tasks by using only task-relevant neurons and prepending adequate task-specific prompt tokens with only a single language model. Experiments reveal that our method significantly enhances inference efficiency (at most ×$\times 1.73$ speed up) for various widely used benchmarks, showing comparable performance to the prompt tuning method. Moreover, our method is applicable across various transformer-based architectures, confirming its practicality and scalability.
Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning
Nakyeong Yang1, Jiwon Moon1, Junseok Kim2, Yunah Jang1 and Kyomin Jung1 1Seoul National University, 2Pohang University of Science and Technology {yny0506, wldnjs913, vn2209, kjung}@snu.ac.kr {junseokkim00}@postech.ac.kr
1 Introduction
With the success of parameter-efficient fine-tuning (PEFT) methods, pre-trained language models perform various tasks by adjusting only small numbers of parameters. Among PEFT methods, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) are widely used to solve multiple tasks since they require only a few tunable tokens to be prepended to the input tokens while freezing the original parameters of a language model (Chan et al., 2023; Huang et al., 2024). This strategy is memory efficient, as it enables a model to solve various tasks with only a few task-specific prompt tokens while retaining the original parameters. However, we can raise one research question regarding this task-specific trait of prompt tuning:
"Leveraging prompt tuning, could we condense task knowledge into a few task neurons in a frozen model to efficiently utilize it?"
Figure 1: Skeleton Framework Overview. Our framework activates only task-localized paths of language models in the prompt tuning procedure. Each color (i.g., red, blue, and yellow) corresponds to a specific task; we obtain a task-specific prompt and neuron indices after executing our framework. Gray-colored paths in the figure represent task-irrelevant paths that are never used during training and inference, thereby improving inference speed.
Existing prompt tuning frameworks are limited by the need to use the full original model architecture during inference. In fact, these methods may require additional inference time, as the prepended prompt tokens introduce further computational overhead. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) have attempted to accelerate the inference speed of prompt-tuned models by pruning unnecessary prompt tokens. However, they have only marginally improved inference efficiency, as the number of pruned tokens is a small fraction ($\leq 0.1\%$) of all the parameters of a language model (Lester et al., 2021).
Prior studies have attempted to investigate and analyze task-relevant neurons in language models to explain the language models’ behavior (Panigrahi et al., 2023; Yang et al., 2024). Specifically, Panigrahi et al. (2023) has suggested training new parameters, which mask the original parameters of language models to determine task neurons. Yang et al. (2024) has utilized an explainability method to determine task-relevant neurons and reveal that some task-relevant neurons significantly affect the outputs of specific tasks (i.e., natural language understanding, social bias). However, while they have introduced a method to identify task-relevant neurons, they fall short of proposing a practical approach to efficiently utilize those neurons during training and inference.
To address this challenge, we propose a novel training framework, Skeleton, designed to enhance the efficiency of language models by leveraging only task-relevant neurons for solving various tasks. Specifically, we use attribution (Yang et al., 2023, 2024), an explainability method, to identify and utilize only task-relevant neurons in the prompt tuning procedure. Our framework requires only a few prompt tokens to be updated and freezes the original parameters of a language model, as it adopts the prompt tuning method. Furthermore, Skeleton dramatically accelerates the inference speed of language models since it utilizes only task-relevant neurons. For example, in order to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only the task-relevant neurons to solve the task.
We demonstrate our method in various widely used natural language understanding benchmarks and reveal that our method significantly outperforms existing PEFT methods in inference efficiency without sacrificing the original performance of each task. Surprisingly, our method improves the inference speed up to ×$\times 1.73$ with 65% of the parameters pruning rate. We additionally conduct various analyses to offer a guideline for using our new prompt tuning framework, such as investigating which types of layers are crucial for processing task knowledge in a language model. Furthermore, our prompt tuning framework is practical and scalable since it is applicable to any transformer-based architecture. In summary, this work makes the following contributions:
  • We first define task neuron and propose Task Neuron Localized Prompt Tuning (Skeleton), a practical training framework for accelerating the inference speed of a language model for various tasks.
  • We reveal that Skeleton successfully enhances the inference speed of each task without compromising the task performance.
2 Backgrounds
2.1 Parameter-efficient Fine-tuning
Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023). Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various NLP tasks (Wang et al., 2023; Bansal et al., 2023). Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model. However, they still face a limitation as they are not parameter-efficient during the inference process. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) propose prompt token selection methods to enhance the inference speed of prompt-tuned language models. Nevertheless, they have only marginally enhanced inference efficiency since the number of pruned tokens is only a tiny proportion of the entire number of input tokens.
2.2 Task Knowledge Localization
In spite of the significant performance of language models, it remains challenging to accurately investigate the role of each parameter in language models in dealing with a specific task. Therefore, existing studies have attempted to identify task-specific neurons in executing a specific task (Panigrahi et al., 2023; Yang et al., 2023, 2024). Panigrahi et al. (2023); Yang et al. (2023, 2024) have studied to localize neurons for performing a specific task by clarifying the role of each parameter in a language model. Panigrahi et al. (2023) has suggested a training-based task knowledge localization method, model grafting, which localizes a task knowledge by training new parameters to mask original parameters. Although it has effectively identified task neurons, it has required excessive additional storage for newly trained masking parameters, which is equivalent to the number of model parameters. Yang et al. (2023, 2024) has illuminated task neurons by utilizing the attribution technique (Shrikumar et al., 2016), an explainability method that derives the importance of each feature when solving a specific task. Yang et al. (2023, 2024) has verified that the attribution effectively detects task neurons for solving a specific task and proposed a task neuron detection method applicable to language modeling tasks. However, Yang et al. (2023, 2024) has a limitation in just pruning the task neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded. Furthermore, the prior study has only proposed the method for pruning task neurons and is far from proposing a practical method to improve the inference speed of task-localized language models. The architecture of transformer variants consists of complex neural computation processes; thus, satisfying the integrity of dimension dependency for adjacent networks while pruning the weights is challenging. We adopt the attribution method (Yang et al., 2023, 2024) and propose a new practical framework for accelerating a language model, where the framework utilizes only task neurons while training a language model via prompt tuning, resolving the challenges in pruning networks in transformer variants.
3 Problem Definition
3.1 Prompt Tuning
We adopt the prompt tuning method to propose a new framework for accelerating a language model. Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)=\prod_{k=1}^{K}\mathcal{P}_{\theta}(y_{k}|x,y_{1},...,y_{k-1})$ that represents a language model, where $\theta$ is parameters of the language model. Prompting is the method of adding extra information to the model to condition during its generation of $y$. Normally, prompting is achieved by adding a series of tokens $P_{\tau}\in\mathbb{R}^{l\times d}$ to the input $x$ as $\mathcal{P}_{\theta}(y|[P_{\tau};x])$, where $l$ and $d$ are token length and feature dimension, and $P_{\tau}$ is trainable parameters for a specific task $\tau$. In the prompt tuning process, we update only $P_{\tau}$ and get $P^{\prime}_{\tau}$ to maximize the likelihood of $y$ using gradient descent while freezing the parameters of the language model $\theta$. After the prompt tuning, we get the new conditional generation function $\mathcal{P}_{\theta}(y|[P^{\prime}_{\tau};x])$.
3.2 Skeleton Framework
In the Skeleton framework, we still update only prompt tokens, freezing the original parameters. However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we first define task neuron as follows:
Task Neuron.
Let $\mathcal{P}_{\theta}$ be a language model, where $\theta=\{\theta_{1},...,\theta_{N}\}$ is a set of parameters of it. Then, $\theta_{\tau}\subset\theta$ is defined as task neurons for a task $\tau$ if the below equation is satisfied:
$\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta}([P^{\prime}_{\tau};x]))\approx\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta_{\tau}}([P^{\prime\prime}_{\tau};x]))$ (1)
where $\mathcal{D}_{\tau}$ means a dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy). $P^{\prime}_{\tau}$ and $P^{\prime\prime}_{\tau}$ are task-specific prompts. $N$ means the total number of parameters in a language model.
Skeleton framework adopts the prompt tuning strategy using only task neurons $\theta_{\tau}$, and get the new task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime\prime}_{\tau};x])$. Therefore, we can utilize the language model more time-efficiently by using only those neurons within the Skeleton framework. In addition to being time efficient, this framework is also memory-efficient since it requires only a small fraction of original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime\prime}_{\tau}$.
4 Methods
In this section, we describe the process of Skeleton. We first train a language model using the prompt tuning method to inject prior knowledge of a specific task into it. After that, we quantify the task relevance of each neuron to determine task neurons in a language model. We finally retrain the prompt tokens with the prompt tuning method to condense the task knowledge to only the task neurons. Note that the original language model maintains only task neurons via a pruning method in this process. After applying Skeleton framework, we can utilize the language model efficiently by using only the task-specific prompt and task neurons, accelerating the inference speed for a specific task. The entire process of Skeleton is described in Figure 2 and Algorithm 1.
Figure 2: Task Neuron Localization Process of Skeleton. After training a model using our framework, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure. Red-colored values in the prune weights, $\tilde{W}_{1}$ and $\tilde{W}_{2}$, mean task neurons for a specific task (task A). The weight of FFN2 is pruned depending on FFN1’s dimension, and the output representation of FFN2 is zero-recovered (Blue-colored values) to match the output dimension of the residual network (described in Section 4.5).
4.1 Quantifying Task Relevance of Neurons
This section describes the process of computing the task relevance of each neuron. We adopt an attribution method (Yang et al., 2023, 2024) to quantify the skill relevance of neurons in the pre-trained language models. Formally, given a task-specific prompt tokens $P_{\tau}$ and text input tokens $x$ provided to $\mathcal{P}_{\theta}$, the contribution of the $i$-th neuron to the prediction of an output $y$ is defined as follows:
$A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$ (2)
where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])/\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron.
We adopt transformer variants; thus, activations and gradients are computed for all input token representations. Therefore, if an input text $x_{j}$ includes a $K_{j}$ number of tokens, each neuron has a $K_{j}$ number of attributions. In addition, there are multiple data instances for each task $\tau$; thus, we aggregate attributions for tokens and instances as follows:
$A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)=\sum_{j}^{N}\sum_{k}^{K_{j}}\frac{1}{K_{j}}A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ (3)
where $A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ means the attribution scores computed for an input text token $t_{k}\in x_{j}$. $\mathcal{D}^{\tau}$ and $N$ mean a task-specific dataset and the number of instances in the dataset, respectively. Therefore, $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ is the attribution score computed for a specific task dataset $\mathcal{D}^{\tau}$. Although $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ can be computed using the entire dataset, we report the experimental results of computing it using only a significantly small amount of data (e.g., only twenty data samples) to ensure the efficiency of our method. To simplify the notation, $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ will be abbreviated as $A_{i}$ from this point forward.
4.2 Identifying Task-relevant Neurons.
After quantifying each neuron’s task relevance, we determine task-relevant neurons and eliminate task-irrelevant neurons using a structured pruning method. We first sort the neurons of the whole target layers by the task relevance scores. Then, we determine the optimal neuron ratio using the Binary search algorithm to retain as few task neurons as possible without compromising the task knowledge. Specifically, we compute the accuracy for the validation set $\mathcal{D}^{val}$ to search for the optimal neuron ratio that does not degrade the task performance. We define a margin $\psi$ and regard $\theta_{\tau}$ as task neurons if the accuracy degradation after neuron localization is smaller than the specified margin $\psi$.
4.3 Task Neuron Localization.
After identifying task neurons, we eliminate task-irrelevant neurons from a language model using a structured pruning method. Suppose that a weight matrix $W\in\mathbb{R}^{d\times m}$ is a linear matrix multiplication parameter, and then the matrix after pruning is denoted as $\tilde{W}=(W_{ij})_{\begin{subarray}{c}1\leq i\leq d\\ j\in\mathcal{M}\end{subarray}}$, where $\mathcal{M}$ is the set of task-relevant neuron indices about the $W$. If the bias term $b\in\mathbb{R}^{m}$ is added to the operation for an affine transformation, the bias term can also be pruned by performing the $\tilde{b}=(b_{i})_{i\in\mathcal{M}}$ operation similarly. The task-localized parameters are used to compute the new representation by performing the transformation operation $h\tilde{W}$ or $h\tilde{W}+\tilde{b}$. Notice that this method is model-agnostic since all neural network models consist of linear transformation layers. For example, transformer variants have self-attention and feed-forward network (FFN) modules, all of which include linear matrix multiplication operations. After localizing a language model, we further intensify task knowledge of a localized language model using a prompt tuning method.
4.4 Task Neuron Localized Prompt Tuning
Our Skeleton framework follows the four steps to train a model: (1) the first training step for injecting prior task knowledge; for injecting prior task knowledge; (2) the quantification step of computing task relevance for each neuron; (3) the task neuron localization step; (4) the final prompt tuning step for condensing task knowledge.
The Prior Knowledge Injection Step.
The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a specific task in order to accurately quantify the task-specific knowledge of each neuron. Therefore, we first execute the prior knowledge injection process. Specifically, we train a language model $\mathcal{P}_{\theta}$ via prompt tuning and get a task-specific prompt $P_{\tau}^{\prime}$.
The Quantification Step.
After injecting task knowledge into the language model, we quantify the task relevance $A_{i}$ of each neuron for $\mathcal{P}_{\theta}$ and $P_{\tau}^{\prime}$ using Equation 2 and 3.
The Task Neuron Localization Step.
After quantifying the task relevance $A_{i}$, we first sort the neurons by the task relevance in descending order. Then, we select and maintain only the top-$p$ neurons $\theta_{\tau}\subset\theta$ from the language model by pruning task-irrelevant neurons (Section 4.3). Note that we use a Binary search algorithm to find the optimal ratio of neurons $p\in[0.0,0.05,...,0.95,1.0]$. This search process requires only a maximum of five trials since we search pruning ratio in 0.05 units.
The Knowledge Condensation Step.
During the task neuron localization step, we condense the task knowledge into only the task neurons. Although we never update the original parameters of the language model, we can condense the task knowledge into the task neurons by updating only $P_{\tau}^{\prime}$. We train the task neuron localized model $P_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$. The Binary search algorithm determines the optimal pruning rate after this step.
Algorithm 1 Skeleton Framework
1:a language model $\mathcal{P}_{\theta}$; a prompt $P_{\tau}$; a margin $\psi$
2:a task neuron localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
3:train $\mathcal{P}_{\theta}$ via prompt tuning using $P_{\tau}$ and get $P^{\prime}_{\tau}$
4:get the validation accuracy $s_{\tau}$ for $\mathcal{P}_{\theta}$ using $P_{\tau}$
5:compute task relevance $A_{i}$ via Equation 2 and 3
6:sort target neurons in descending order by $A_{i}$
7:
8:$l\leftarrow 0$
9:$h\leftarrow 20$
10:while $l\leq r$ do                         // Binary search
11:     $m\leftarrow\lfloor(l+h)/2\rfloor$
12:     $p=m\times 0.05$
13:     get top-$p$ task neuron localized model $\mathcal{P}_{\theta_{\tau}}$
14:     train $\mathcal{P}_{\theta_{\tau}}$ via prompt tuning using $P^{\prime}_{\tau}$ and get $P^{\prime\prime}_{\tau}$
15:     get validation accuracy $s^{\prime}_{\tau}$ for $\mathcal{P}_{\theta_{\tau}}$ using $P^{\prime\prime}_{\tau}$
16:     if $s_{\tau}-s^{\prime}_{\tau}\leq\psi$ then
17:         $\mathcal{P}_{\theta_{\tau}}^{*}\leftarrow\mathcal{P}_{\theta_{\tau}}$             // The optimal model
18:         $l\leftarrow m+1$
19:     else
20:         $h\leftarrow m-1$
21:return $\mathcal{P}_{\theta_{\tau}}^{*}$
Model Method SST-2 IMDB AGNews MRPC CB
BERT (110M) Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57
Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76
Prompt-tuning 89.36 88.56 90.58 67.07 69.63
Skeleton 89.21 88.65 89.94 68.05 69.04
[…]
RoBERTa (125M) Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90
Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02
Prompt-tuning 93.84 92.72 91.87 67.23 66.06
Skeleton 92.81 91.50 90.79 67.89 69.04
[…]
Table 1: Task Neuron Localized Prompt Tuning Results. We report the accuracy of five datasets for Skeleton and other baselines. We train models using each method by three trials and report the averaged accuracy. Pruning Rate (%) and Speed up ($\times$) mean the ratio of neurons deactivated after applying our method and the inference speed up compared to the Prompt-tuning, respectively. The values in parenthesis are the number of pruned parameters. The pruning rates are calculated based on the maximum number of prunable parameters.
4.5 Details of the Skeleton Framework
Satisfying Dimension Integrity.
This section describes the detailed task localization processes; the entire process is shown in Figure 2. We prune task-irrelevant neurons from a language model. Recent language models have been trained using the transformer architecture, but they consist of complex neural computation processes. For example, sequentially connected networks are dependent on each other in terms of dimensionality. Suppose there are two sequential feed-forward networks (FFNs), $W_{1}\in\mathbb{R}^{d\times m}$ and $W_{2}\in\mathbb{R}^{m\times d}$. If we prune the first FFN and get a pruned weight $\tilde{W}_{1}\in\mathbb{R}^{d\times m^{\prime}}$, this pruning process affects the inconsistency of the dimension with the weight in the second FFN. Specifically, $W_{2}$ is forced to be pruned as $\tilde{W}_{2}\in\mathbb{R}^{m^{\prime}\times d}$. Therefore, we implement the task neuron localization algorithm by reflecting the dimension dependency of two FFNs. In addition, suppose the representation computed by the second FFN, $v\in\mathbb{R}^{d}$, is connected to the representation of a residual connection network, $u\in\mathbb{R}^{d}$. Then, if we prune the second FFN and get a pruned weight $\tilde{W}_{2}\in\mathbb{R}^{m\times d^{\prime}}$, this process also triggers the inconsistency of the dimension with the representation transferred by the residual connection, making $v\in\mathbb{R}^{d^{\prime}}$. Thus, we insert zero values into the pruned dimension of the second FFN’s representation, $v$, to ensure its integrity.
Accelerating Inference Speed.
Our framework does not just convert the pruned value of weights to zero values but eliminates the entire column (i.e., neuron) of them. Therefore, our framework accelerates the language model inference speed. We do not consider other types of networks (e.g., Attention modules, Embedding layer, Language model head) since they play a crucial role in the model’s language understanding.
Furthermore, the parameters of FFNs constitute most of the prunable parameters (about more than $65\%$); thus, pruning even only the FFNs is sufficient. The experimental results for other modules are described in Section 5.4.
5 Experiments
5.1 Experimental Setup
Datasets.
We conduct experiments on five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018). Specifically, we utilize SST-2, IMDB (sentiment analysis); AGNews (topic classification); MRPC (semantic textual matching); CB (natural language inference) to verify the applicability of our method. We split the train set and use 10% of them as a validation set.
Models and Training Methods.
We utilize two transformer variants  [^2] https://huggingface.co/docs/transformers/, BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), for our experiments since they are some of the most popular basic language models. We select a prompt tuning method (Lester et al., 2021) to compare with our method, and also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
Implementation Details.
We evaluate Skeleton and baselines on NVIDIA A5000 GPU. We train models using $l=20$ prompt tokens during 100 epochs with early-stop conditions for all datasets. In addition, we set the margin $\psi=1.0\%$ for the validation accuracy score used in the Binary search algorithm. We search $p\in[0.0,0.05,...,0.95,1.0]$ to find the best pruning rate in the Binary search algorithm. We quantify the task relevance using only $n=20$ data instances of each training dataset, following Yang et al. (2024). We specify the target modules for task neuron localization as FFN networks since the experimental results for FFN networks outperform those of other modules, shown in Figure 6. We use learning rates $\gamma_{1}\in$ [1e-5, 5e-5] and $\gamma_{2}\in$ [1e-4, 5e-3] for fine-tuning and prompt-tuning, respectively. We use a masked language model probing network including an instruction for solving a task to utilize the pre-trained knowledge of BERT and RoBERTa, as shown in Appendix A.
5.2 Skeleton is Efficient and Robust
Datasets Experiments.
We evaluate and report the task performance and efficiency of each method in Table 1 for the five natural language understanding benchmarks. We adopt BERT-base and RoBERTa-base to conduct the experiments. The results show that our method performs comparably to the prompt tuning method and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% of pruning rates). Surprisingly, the pruned networks show at most ×$\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models. We describe the detailed elapsed time (sec) of Skeleton models and prompt tuned model in Figure 3. These results reveal that our method significantly outperforms the prompt tuning method in inference speed.
Figure 3: Elapsed Time and Inference Speed-up. We plot the elapsed time and inference speed-up for BERT-base and RoBERTa-base on the four datasets. The values above the bar plot are the speed-up ratios.
Hyperparameter Experiments.
We conduct experiments to investigate the time efficiency of our framework for varying hyperparameters, as shown in Figure 4. Specifically, we search varying batch sizes and the number of input tokens to demonstrate efficiency by comparing our framework to the prompt tuning method. We sample random texts with $k$ tokens and input them into the BERT-base model for the experiments. For the batch size experiments, we evaluate the inference speed with varying batch sizes $\beta\in[16,32,64,256,512]$, fixing the number of tokens at $64$. For the number of tokens experiments, we measure the inference speed with multiple numbers of tokens $k\in[32,64,256,512]$, using the batch size of $64$. The results reveal that our framework shows robust efficiency improvement in various hyperparameters, showing better speed in higher batch sizes and a widely used number of tokens.
Figure 4: Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
5.3 Task Localization Methods Experiments
We compare our method to other task neuron selection methods: Activation, Gradient, and Random, to justify the excellence of the attribution method (Section 5.3). Activation identifies task neurons by using the activated values of neurons. We compute the sum of activated values of each neuron when inputting a dataset. Gradient selects task neurons by using the gradient values of each neuron. We compute the sum of the gradient values of each neuron when predicting the label of the utilized dataset. Random randomly selects task neurons.
We demonstrate the task localization performance of those baselines to justify our selection of the task localization method, Attribution, as shown in Figure 5. Specifically, we select the BERT-base model for the SST-2 and IMDB datasets for the experiments. We prune the top 50% of task neurons by using those methods for a fair comparison. Surprisingly, even Activation and Gradient do not show competitive performance compared to Random. However, Skeleton (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
Figure 5: Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
5.4 Module-specific Task-localization
Transformer variants have various types of modules (i.g., attention modules and FFNs). Therefore, we examine which module is suitable for maintaining the knowledge of a language model (BERT-base). Specifically, we categorize layers into four segments: (1) All layers (All), (2) Attention layers (Attn), (3) Dense layers (Dense), and (4) Feed-forward layers (FFN). All includes all linear layers. Attn consists of Query (Q), Key (K), Value (V), and Output (O) layers of the attention module. Dense contains $O$ of the attention module and two FFNs. FFN includes two FFNs. We prune the top 30% of task neurons from each module for a fair comparison. Figure 6 represents the results of our method for various types of modules. From these experiments, we reveal that FFNs are suitable for maintaining the knowledge of a language model in our prompt tuning setting. Adopting the FFN module yields faster performance than other modules without compromising accuracy.
Figure 6: Module-specific Task-localization Results. We plot the relationship between the accuracy and speed up of each module, where each diagram corresponds to the categorized module. The red-dotted line is the baseline performance (Prompt-tuning).
5.5 Prior Knowledge Injection is Important
Figure 7: Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
This section describes the importance of injecting prior task knowledge (described Section 4.4) to accurately quantify the task relevance, as shown in Figure 7. We prune the top 50% of task neurons by using two methods (i.e., Skeleton and w/o Prior Task Knowledge) for a fair comparison. The experimental results reveal that prior task knowledge injection has a crucial role in quantifying the task relevance. Excluding the Prior Knowledge Injection step results in slower convergence and lower task accuracy compared to our method.
5.6 Skeleton is Scalable to a Larger Model
We conduct experiments to verify that our method is scalable to a larger model, as shown in Table 2. Specifically, we adopt BERT-large to show the scalability of our framework. These results reveal that our method is scalable to a larger model.
Method SST-2 IMDB
BERT-base (110M) Fine-tuning (Single) 92.39 92.50
Fine-tuning (Multi) 91.36 92.16
Prompt-tuning 89.36 88.56
Skeleton 89.21 88.65
[…]
BERT-large (340M) Fine-tuning (Single) 92.77 93.57
Fine-tuning (Multi) 92.73 93.83
Prompt-tuning 91.28 89.96
Skeleton 92.19 89.35
[…]
Table 2: Experiments on various sizes of BERT.
6 Conclusion
In this study, we propose Skeleton, which explores the extreme of the prompt tuning method by localizing task-relevant neurons from the original language model. Skeleton investigates the optimal task-localized sub-network from a language model and condenses task knowledge by utilizing only task-relevant parameters during the prompt-tuning procedure. We demonstrate our method for various benchmarks and reveal that Skeleton dramatically enhances the efficiency of language models, showing significant inference speed boost-up.
Limitations
Our framework accelerates the inference speed of language models, identifying the optimal subnetwork for utilizing language models. However, while our approach exhibited comparable performance and occasionally slightly superior results to the existing prompt tuning method, these improvements were not consistent across overall task performance. In addition, our experiments are limited to five widely-used natural language understanding datasets; thus, we should extend to a wide range of datasets to validate the applicability of our framework. Furthermore, we select feed-forward networks to localize task neurons in language models since they are adequate for inference speed acceleration. However, the entire modules in Transformer should be considered as a localization target to achieve the extreme inference speed up; we leave this point as future research.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea.
Appendix A Instruction Templates.
We use text templates for input texts of the utilized datasets. Specifically, the text templates consist of an instruction and an input text, as shown in Table 3. For the prediction, we utilize the representation of the mask token (i.e., "<mask>") to utilize the pre-trained knowledge of encoder-only language models.
Appendix B Module-specific Task-localization Details
We conduct experiments for various modules in the transformer architecture, as described in Section 5.4. We implement algorithms for pruning those networks following our Skeleton framework to satisfy the dimension of adjacent networks after pruning. For example, suppose that we prune the Key network, $W^{K}\in\mathbb{R}^{d\times d}$, and get a pruned weight, $\tilde{W}^{K}\in\mathbb{R}^{d\times d^{\prime}}$, in the Attention module. Then, we should also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since they are adjacent networks. Therefore, after the pruning process, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by adding zero values to ensure the identical dimension. We also implement other layers in the same way with it.